OpenStudy (anonymous):

what is the uncertainty in the location of a photon of wavelength 5000 Angstrom if wavelength is known to an accuracy of 1 pm?

OpenStudy (jfraser):

use the uncertainty equation \[\Delta x * \Delta mv \ge\frac{h}{4 \pi}\]

OpenStudy (jfraser):

and solve for \(\Delta x\)

OpenStudy (anonymous):

oh! thnx 4 d help!! :)

OpenStudy (anonymous):

thnx